the perimeter of an isosceles triangle is 49m. if the base is twice the side, what is the length of the side?
Answer by stanbon(75887)

the perimeter of an isosceles triangle is 49m. if the base is twice the side, what is the length of the side?
----
Let each of the equal sides be "x".
Then base 2x.
----
Equation:
x + x + 2x = 49m
---
4x = 49m
x = (49/4)meters = 12 1/4 meters (length of each of the equal sides)
===========
Cheers,
Stan H.